OpenStudy (anonymous):

Having difficulty graphing... When rabbits were first brought to Australia, they increased rapidly in numbers. Assume that in 1861, there were 38 rabbits, and in 1863 there were 1500 rabbits, in 1864 there were 9487 rabbits, and in 1865 there were 60,000 rabbits. Use 61 for 1861, 63, for 1863, etc., or your calculator will give you an error. Make a scatter plot of this data. In 1870, according to your model, how many rabbits would there be?

OpenStudy (anonymous):

a) 27,357,822 b) 7,293,462,308 c) 973,724 d) 595,334,768

OpenStudy (anonymous):

@jim_thompson5910 Please help!

OpenStudy (anonymous):

sry not quite sure on this one

OpenStudy (anonymous):

@chantz417 It's alright! I was able to graph it but I do not know how to find 1870!

OpenStudy (anonymous):

@lshiny give me a minute I think I can help.

OpenStudy (anonymous):

@GLKK Okay, thank you!

OpenStudy (anonymous):

@lshiny could you show me your graph?

OpenStudy (anonymous):

I don't know how I am able to show you :( I'm sorry. I just punched the options in my graphing calculator.

OpenStudy (anonymous):

@lshiny alright. give me a minute.

OpenStudy (anonymous):

I would go with C. Seems the most reasonable. I could be wrong though.

OpenStudy (anonymous):

@lshiny

OpenStudy (anonymous):

i was thinking A

OpenStudy (anonymous):

because in one year it went up 49,000 and there is 7 years ahead that you need to find

jimthompson5910 (jim_thompson5910):

This problem is just really unrealistic. It assumes exponential growth which accelerates very very quickly. It doesn't take into account any other factors like predators or lack of food to slow the population growth rate down.

OpenStudy (anonymous):

@jim_thompson5910 agreed. I also misread. I think C or A would be your best answers. @lshiny

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Anyways, what you do is plot all of the points on a scatter plot like this

OpenStudy (anonymous):

I agree completely. I am leaning towards C due to the fact that the past year gaps didn't have such a dramatic growth...

OpenStudy (anonymous):

@jim_thompson5910 What do you think?

OpenStudy (anonymous):

@lshiny I would go for C.

OpenStudy (anonymous):

^ Yeah I'm learning towards C too! Thanks @GLKK for helping!

jimthompson5910 (jim_thompson5910):

then you use exponential regression on a calculator to get this exponential function y = (6.39967*10^(-48)) * 6.303060573^x and then you plug in x = 70 to get y = (6.39967*10^(-48)) * 6.303060573^x y = (6.39967*10^(-48)) * 6.303060573^70 y = 595,334,749.644288 due to rounding errors, it's going to be a bit off of what they want, which is 595,334,768

jimthompson5910 (jim_thompson5910):

so you can see why this model is ridiculous, especially in a very brutal environment such as Australia

OpenStudy (anonymous):

Whoa, @jim_thompson5910 That was completely correct. I totally forget that this is possible through exponential regression. Thank you so much!

jimthompson5910 (jim_thompson5910):

this is what geogebra gives me, which is the most accurate tool I could find so far that red curve does its best job to get as close as possible to each point then I computed f(70) to be 595,334,768.3